# Realistic models of general-relativistic differentially rotating stars

### v1

@@ 1 Introduction
While these laws have been extensively studied in the literature and applied to the analysis of differentially rotating neutron stars and quark stars, they do not always represent a good match to the simulations. To improve the level of realism in the construction of equilibrium models of differentially rotating neutrons stars, we have here reconsidered this problem with a bottom-up approach in which we first consider in detail the rotational information obtained from three, state-of-the-art BNS merger simulations performed with different and realistic EOSs and mass ratios. From these simulations we have then extracted the quasi-stationary rotational properties of the remnants, both in terms of azimuthally- and time-averaged profiles of the angular velocity and of the specific angular momentum. In this way, it was possible to study the existence of universal features in the angular-velocity profile and to point out a criterion that allows us to simply recognise the transition between the “core” of the remnant and the portion of the HMNS whose angular-velocity profile follows a Keplerian behaviour, i.e., the “disc”.
@@ 2.1 Numerical setup
A well-known criterion, the “turning-point” criterion (Friedman et al., 1988) then states that the condition (15) is a sufficient condition for the quasi-radial dynamical instability leading to a gravitational collapse to black holes. A less well-known result is that the “turning-point stability line”, i.e., the line marked by all stellar models satisfying Eq. (15), is distinct (for $J>0$) from the “neutral-stability line”, which is instead marked by all stellar models whose fundamental $f$-mode frequency vanishes (Takami et al., 2011). By construction, and independently of the type of rotation law, such a line represents a necessary condition for the appearance of a dynamical instability leading to a gravitational collapse, hence, truly distinguishing stellar models that are unstable from those that are instead stable. Unfortunately, the neutral-stability line can be determined either via complex perturbative studies (Dimmelmeier et al., 2006; Zink et al., 2010; Takami et al., 2011) or through direct numerical-relativity calculations (Font et al., 2002; Takami et al., 2011; Weih et al., 2018; Szewczyk et al., 2023). Because of its simplicity, and bearing in mind that the differences between the two lines in terms of masses and central rest-mass densities are of a few percent at most, we will hereafter make use only of the turning-point stability line.
@@ 3.1.2 Varying $A,B,p$ and $q$
Figure 3: Angular velocity as a function of the specific angular momentum $j$ for a stellar model obeying the DD2 EOS, having a central rest-mass density $\rho_{c}=8.024\times 10^{14}\,{\rm g~{}cm^{-3}}$, and an axis ratio $R_{p}/R_{e}=0.8$. The underlying differential rotation law is U8 and the different panels show how the functional form changes when varying the relevant parameters $\hat{A},\hat{B},p,q$ indicated in the top-right part of each panel. Note that each panel contains an inset reporting, for the same models, the angular velocity as a function of the proper circumferential radial coordinate $R$.
@@ 3.2 The Uryu-9 rotation law
$\lambda_{2}:=\frac{\Omega_{e}}{\Omega_{c}}=\left(1+\left(\frac{j_{e}}{B^{2}\Omega_{c}}\right)^{p}\right)\left(1-\frac{j_{e}}{A^{2}\Omega_{c}}\right)\,.$ (35)
@@
After a bit of algebra, it is possible to show that the expressions for $A$ and $B$ appearing in (34) and (35) follow from the solution of the quadratic equation in $B$
@@
We note that the numerical calculation of equilibrium models with the U9 law can lead to numerical challenges, at least with the Hydro-RNS code. In particular, convergence often fails at large values of $\lambda_{2}$, where $\Omega_{e}>\Omega_{c}$, as the iterated equatorial radius $r_{e}$ alternates between two different solutions with different masses and radii for the same value of the axis ratio (these two branches merge again at small axis ratios). Considering that the branch yielding lower masses also exhibits an unphysical relation between the mass $M$ and the specific angular momentum $j$ (i.e., $M(j)$ is no longer injective with $j$) and non-smooth profiles in $\Omega(r)$ as well as $M(j)$, we consider such branch as “unphysical” and discuss hereafter only the results relative to the “physical” branch, that is, the branch yielding larger masses for the same axis ratio.
@@ 4 Differential-rotation laws from BNS merger simulations
When building differentially rotating equilibria that are meant to model the HMNS produced in the merger of a binary system of neutron stars, it is clear that the only way to obtain realistic differential-rotation profiles it to extract this information from accurate and general-relativistic simulations of binary mergers using a state-of-the-art microphysical description. To this scope, we consider the results of BNS-merger simulations with three different temperature-dependent EOSs: DD2 (Hempel & Schaffner-Bielich, 2010b), TNTYST (Togashi et al., 2017), and the V-QCD of intermediate stiffness (Demircik et al., 2020; Demircik et al., 2022).
@@
More precisely, the first remnant we consider is the product of an equal-mass BNS with component masses $M_{1,2}=1.4\,M_{\odot}$ evolved with the DD2 EOS by Ng et al. (2024); since it survives for times $t-t_{\rm mer}\gtrsim 100\,{\rm ms}$ it represents a very good candidate to study how the properties and rotation profiles change with time and reach a quasi-stationary equilibrium. The second simulation considers the merger of a magnetized equal-mass binary with component masses $M_{1,2}=1.275\,M_{\odot}$ and the TNTYST EOS (Togashi et al., 2017). In this case, a dipolar magnetic field is present inside the star with maximum magnetic-field strength $B\sim 10^{14}\,{\rm G}$ (Chabanov et al., 2023) and the merger remnant is an HMNS that does not collapse promptly to a black hole. Finally, the third merger remnant refers to an unequal-mass binary with mass ratio $q:=M_{2}/M_{1}=0.7$, where $M_{1}=1.64\,M_{\odot}$ and $M_{2}=1.15\,M_{\odot}$. The binary is evolved with the V-QCD EOS of intermediate stiffness, where the temperature dependence is introduced at low densities via the HS-DD2 EOS (including relativistic mean-field interactions) (Hempel & Schaffner-Bielich, 2010a; Typel et al., 2010) and in the dense nuclear-matter phase via a van der Waals model (Demircik et al., 2022). In this case, the merger remnant has a lifetime of $t-t_{\rm mer}\gtrsim 37\,{\rm ms}$, which, as we will show below, is sufficient to extract robust rotation profiles. The basic properties of the binaries considered here and some of the corresponding dynamical properties have been summarised in Table 1.
@@
Table 1: Properties of the BNS merger configurations. Reported are the initial masses $M_{1}$, $M_{2}$, the EOS, and the collapse time $t_{\rm coll}$. Also reported are the coefficients of the linear fit (47), the time $t-t_{\rm mer}$ at which is performed, and the corresponding values for the core radius $R_{\rm core}$, the rest-mass density $\rho_{\rm core}$ and angular velocity $\Omega_{\rm core}$ of the core.
EOS $M_{1}$ $M_{2}$ $t_{\rm coll}$ $t-t_{\rm mer}$ $a\times 10^{-2}$ $b$ $\rho_{\rm core}\times 10^{-13}$ $\Omega_{\rm core}/2\pi$ $R_{\rm core}$
@@ 4.2 Quasi-stationarity and rotation profiles
For each of the simulations discussed above, we analyse the post-merger rotation profile of the long-living remnants, extracting rotation quantities such as the angular velocity $\Omega$ and the specific angular momentum $j$. Furthermore, since the merger remnant has an equilibrium that varies rapidly right after the merger, attention needs to be paid to make sure that the rotational properties are measured when the remnant has reached a quasi-stationary equilibrium. To this scope, we define a time window after the merger with extremes $t-t_{\rm mer}=t_{1}$ and $t-t_{\rm mer}=t_{2}$ (where $0<t_{1}<t_{2}$) during which we monitor the evolution of properties of the merger remnant. In particular, we use $t_{1}$ as a reference time from when the system has started to reach a quasi axisymmetric equilibrium and $t_{2}$ to mark the time when this equilibrium has been reached to a given tolerance.
@@ 4.3 Distinguishing core and disc in the remnant
A few remarks are worth making. First, note that the point of departure from a linear-in-log behaviour of the angular velocity with the rest-mass density coincides with the corresponding departure of the angular velocity from a Keplerian profile with radius. This is not surprising but underlines that the point of departure can be interpreted either as the “beginning” of the Keplerian disc, i.e., $R_{\rm core}$ or as the “end” location of where the density in the HMNS stops growing as a power-law with radius, i.e., $\rho_{\rm core}$. Second, note that the “core-disc” transition does not coincide with the maximum of the angular velocity, which systematically lies at smaller radii; this point, which was already remarked by Hanauske et al. (2017), warns against using the angular velocity maximum as a particularly significant location in the HMNS. Third, note that at least in the limited sample considered here $\rho_{\rm core}$ varies by a factor two at most, i.e., between $\simeq 3\times 10^{13}\,{\rm g/cm}^{3}$ and $\simeq 3\times 10^{13}\,{\rm g/cm}^{3}$ and $\simeq 6\times 10^{13}\,{\rm g/cm}^{3}$ (see Table 1). Finally, note that $R_{\rm core}$ has a much smaller variance (i.e., $\lesssim 7\%$) and while a more systematic analysis may constrain this value more precisely, it is interesting that very different binaries lead to essentially the same core radius.
All things considered, the suggestion of determining the start of the disc as the location where the angular velocity does no longer follow a power-law with the rest-mass density appears to be easy to employ and physically well-grounded; furthermore it points out that values $\rho_{\rm core}\simeq 4\times 10^{13}\,{\rm g/cm}^{3}$ and $R_{\rm core}\simeq 12\,{\rm km}$ can be taken as a reasonable references for the core density and radius, respectively.
@@ 5.2 A new differential-rotation law
As we will discuss in more detail in Sec. 5.3, the added flexibility introduced in the U9+ law allows it to provide a rather accurate description of realistic differential-rotation profiles measured from BNS mergers. At the same time, it can be further improved via a novel differential-rotation law that is directly informed from the simulations and there the angular velocity can be modeled as
@@
Note that when implementing the rotation law (55) in the Hydro-RNS code we need to introduce the standard (rescaled) parameters $\hat{j}_{\rm max},\hat{j}_{e},\lambda_{1},\lambda_{2}$, and $A$, but also the new parameters $\beta$ and $\gamma$. We postpone to Sec. 5.3 the discussion of how the new differential-rotation law – that hereafter will refer to as the “CR” law – performs when modelling the remnants of BNS mergers, we here briefly provide evidence that this law allows one to construct numerical models of differentially rotating stars by presenting in Fig. 11 a representative space of solutions for stars modelled with the DD2 EOS and with the following set of fixed parameters: $\hat{j}_{\rm max}/\hat{j}_{e}=0.82$, $\beta=0.95$, $\gamma=1.8$, $A=20$, $\lambda_{1}=3$ and that was obtained after performing a progressive linear blending with the U9+ law, as discussed in Appendix B. Note that also in this case, the choice of parameters leads to a turning-point stability line such that $\partial M/\partial\rho_{c}|_{\rm tp}>0$. As we comment next, this is not a coincidence and reflects a general behaviour of differential-rotation laws that are inspired by BNS mergers.

### v2

@@ 1 Introduction
These laws have been extensively studied in the literature and applied to the analysis of differentially rotating neutron stars and quark stars. While they already represent a good match to the simulations, they be further refined and extended. In fact, to increase the level of realism in the construction of equilibrium models of differentially rotating neutrons stars, we have here reconsidered this problem with a bottom-up approach in which we first consider in detail the rotational information obtained from three, state-of-the-art BNS merger simulations performed with different and realistic EOSs and mass ratios. From these simulations we have then extracted the quasi-stationary rotational properties of the remnants, both in terms of azimuthally- and time-averaged profiles of the angular velocity and of the specific angular momentum. In this way, it was possible to study the existence of universal features in the angular-velocity profile and to point out a criterion that allows us to simply recognise the transition between the “core” of the remnant and the portion of the HMNS whose angular-velocity profile follows a Keplerian behaviour, i.e., the “disc”.
@@ 2.1 Numerical setup
A well-known criterion, the “turning-point” criterion (Friedman et al., 1988) then states that the condition (15) is a sufficient condition for the quasi-radial dynamical instability leading to a gravitational collapse to black holes. A less well-known result is that the “turning-point stability line”, i.e., the line marked by all stellar models satisfying Eq. (15), is distinct (for $J>0$) from the “neutral-stability line”, which is instead marked by all stellar models whose fundamental $f$-mode frequency of quasi-radial oscillations vanishes (Takami et al., 2011). By construction, and independently of the type of rotation law, such a line represents a necessary condition for the appearance of a dynamical instability leading to a gravitational collapse, hence, truly distinguishing stellar models that are unstable from those that are instead stable. Unfortunately, the neutral-stability line can be determined either via complex perturbative studies (Dimmelmeier et al., 2006; Zink et al., 2010; Takami et al., 2011) or through direct numerical-relativity calculations. More specifically, time-dependent simulations in full general relativity can be carried out by perturbing rotating and nonrotating stellar models and by measuring the corresponding $f$-mode frequencies, which become increasingly smaller as the stellar models approach the neutral-stability line. A multi-dimensional fitting procedure is then applied to extrapolate such frequencies to vanishing values, hence obtaining the neutral-stability line in the $(\rho_{c},M$) plane.
calculations Because of its simplicity, and bearing in mind that the differences between the two lines in terms of masses and central rest-mass densities are of a few percent at most, we will hereafter make use only of the turning-point stability line.
@@ 3.1.2 Varying $A,B,p$ and $q$
Figure 3: Angular velocity as a function of the specific angular momentum $j$ for a stellar model obeying the DD2 EOS, having a central rest-mass density $\rho_{c}=8.024\times 10^{14}\,{\rm g~{}cm^{-3}}$, and an axis ratio $r_{p}/r_{e}=0.8$. The underlying differential rotation law is U8 and the different panels show how the functional form changes when varying the relevant parameters $\hat{A},\hat{B},p,q$ indicated in the top-right part of each panel. Note that each panel contains an inset reporting, for the same models, the angular velocity as a function of the proper circumferential radial coordinate $R$. In addition, in the top row the parameters $\hat{A}$ and $\hat{B}$ are varied but such that their ratio remains constant, thus leading to three different curves only.
@@ 3.2 The Uryu-9 rotation law
$\lambda_{2}:=\frac{\Omega_{e}}{\Omega_{c}}=\left(1+\left(\frac{j_{e}}{B^{2}\Omega_{c}}\right)^{p}\right)\left(1-\frac{j_{e}}{A^{2}\Omega_{c}}\right)\,.$
@@
After a bit of algebra, it is possible to show that the expressions for $A$ and $B$ appearing in (35) and (35) follow from the solution of the quadratic equation in $B^{2}$
@@
We note that the numerical calculation of equilibrium models with the U9 law can become problematic and convergence be lost, at least with the Hydro-RNS code. In particular, convergence often fails at large values of $\lambda_{2}$, where $\Omega_{e}>\Omega_{c}$, as the iterated equatorial radius $r_{e}$ alternates between two different solutions with different masses and radii for the same value of the axis ratio (these two branches merge again at small axis ratios). Considering that the branch yielding lower masses also exhibits an unphysical relation between the mass $M$ and the specific angular momentum $j$ (i.e., $M(j)$ is no longer injective with $j$) and non-smooth profiles in $\Omega(r)$ as well as $M(j)$, we consider such branch as “unphysical” and discuss hereafter only the results relative to the “physical” branch, that is, the branch yielding larger masses for the same axis ratio.
@@ 4 Differential-rotation laws from BNS merger simulations
When building differentially rotating equilibria that are meant to model the HMNS produced in the merger of a binary system of neutron stars, it is clear that the only way to obtain realistic differential-rotation profiles it to extract this information from accurate and general-relativistic simulations of binary mergers using a state-of-the-art microphysical description. Here, we consider the results of BNS-merger simulations with three different temperature-dependent EOSs: DD2 (Hempel & Schaffner-Bielich, 2010b), TNTYST (Togashi et al., 2017), and the V-QCD of intermediate stiffness (Demircik et al., 2020; Demircik et al., 2022).
@@
More precisely, the first remnant we consider is the product of an equal-mass BNS with component masses $M_{1,2}=1.4\,M_{\odot}$ evolved with the DD2 EOS by Ng et al. (2024); since it survives for times $t-t_{\rm mer}\gtrsim 100\,{\rm ms}$ it represents a very good candidate to study how the properties and rotation profiles change with time and reach a quasi-stationary equilibrium. The second simulation considers the merger of a magnetized equal-mass binary with component masses $M_{1,2}=1.275\,M_{\odot}$ and the TNTYST EOS (Togashi et al., 2017). In this case, a dipolar magnetic field is present inside the star with maximum magnetic-field strength $B\sim 10^{14}\,{\rm G}$ (Chabanov et al., 2023) and the merger remnant is an HMNS that does not collapse promptly to a black hole. In practice, even the presence of an initial magnetic field of $10^{14}\,{\rm G}$ does not modify the dynamics of the remnant significantly and we observe a behaviour in terms of differential rotation that is very similar to the case of zero magnetic fields. It may be necessary to consider much larger magnetic-field strengths, e.g., $10^{14}-10^{17}\,{\rm G}$ to appreciate differences in the dynamics and equilibria of the remnant. Finally, the third merger remnant refers to an unequal-mass binary with mass ratio $q:=M_{2}/M_{1}=0.7$, where $M_{1}=1.64\,M_{\odot}$ and $M_{2}=1.15\,M_{\odot}$. The binary is evolved with the V-QCD EOS of intermediate stiffness, where the temperature dependence is introduced at low densities via the HS-DD2 EOS (including relativistic mean-field interactions) (Hempel & Schaffner-Bielich, 2010a; Typel et al., 2010) and in the dense nuclear-matter phase via a van der Waals model (Demircik et al., 2022). In this case, the merger remnant has a lifetime of $t-t_{\rm mer}\gtrsim 37\,{\rm ms}$, which, as we will show below, is sufficient to extract robust rotation profiles. The basic properties of the binaries considered here and some of the corresponding dynamical properties have been summarised in Table 1.
@@
Table 1: Properties of the BNS merger configurations. Reported are the initial masses $M_{1}$, $M_{2}$, the EOS, and the collapse time $t_{\rm coll}$. Also reported are the coefficients of the linear fit (47), the time $t-t_{\rm mer}$ at which is performed, and the corresponding values for the core radius $R_{\rm core}$, the rest-mass density $\rho_{\rm core-disc}$ and angular velocity $\Omega_{\rm core}$ of the core.
EOS $M_{1}$ $M_{2}$ $t_{\rm coll}$ $t-t_{\rm mer}$ $a\times 10^{-2}$ $b$ $\rho_{\rm core-disc}\times 10^{-13}$ $\Omega_{\rm core}/2\pi$ $R_{\rm core}$
@@ 4.2 Quasi-stationarity and rotation profiles
For each of the simulations discussed above, we analyse the post-merger rotation profile of the long-living remnants, extracting rotation quantities such as the angular velocity $\Omega$ and the specific angular momentum $j$. Furthermore, since the merger remnant has an equilibrium that varies rapidly right after the merger, attention needs to be paid to make sure that the rotational properties are measured when the remnant has reached a quasi-stationary equilibrium. Thus, we define a time window after the merger with extremes $t-t_{\rm mer}=t_{1}$ and $t-t_{\rm mer}=t_{2}$ (where $0<t_{1}<t_{2}$) during which we monitor the evolution of properties of the merger remnant. In particular, we use $t_{1}$ as a reference time from when the system has started to reach a quasi axisymmetric equilibrium and $t_{2}$ to mark the time when this equilibrium has been reached to a given tolerance.
@@ 4.3 Distinguishing core and disc in the remnant
A few remarks are worth making. First, note that the point of departure from a linear-in-log behaviour of the angular velocity with the rest-mass density coincides with the corresponding departure of the angular velocity from a Keplerian profile with radius. This is not surprising but underlines that the point of departure can be interpreted either as the “beginning” of the Keplerian disc, i.e., $R_{\rm core}$ or as the “end” location of where the density in the HMNS stops growing as a power-law with radius, i.e., $\rho_{\rm core-disc}$. Second, note that the “core-disc” transition does not coincide with the maximum of the angular velocity, which systematically lies at smaller radii; this point, which was already remarked by Hanauske et al. (2017), warns against using the angular velocity maximum as a particularly significant location in the HMNS. Third, note that at least in the limited sample considered here $\rho_{\rm core-disc}$ varies by a factor two at most, i.e., between $\simeq 3\times 10^{13}\,{\rm g/cm}^{3}$ and $\simeq 3\times 10^{13}\,{\rm g/cm}^{3}$ and $\simeq 6\times 10^{13}\,{\rm g/cm}^{3}$ (see Table 1). Finally, note that $R_{\rm core}$ has a much smaller variance (i.e., $\lesssim 7\%$) and while a more systematic analysis may constrain this value more precisely, it is interesting that very different binaries lead to essentially the same core radius.
All things considered, the suggestion of determining the start of the disc as the location where the angular velocity does no longer follow a power-law with the rest-mass density appears to be easy to employ and physically well-grounded; furthermore it points out that values $\rho_{\rm core-disc}\simeq 4\times 10^{13}\,{\rm g/cm}^{3}$ and $R_{\rm core}\simeq 12\,{\rm km}$ can be taken as a reasonable references for the core density and radius, respectively.
@@ 5.2 A new differential-rotation law
As we will discuss in more detail in Sec. 5.3, the added flexibility introduced in the U9+ law allows it to provide a rather accurate description of realistic differential-rotation profiles measured from BNS mergers. At the same time, it can be further improved via a novel differential-rotation law that is directly informed from the simulations and there the angular velocity can be modelled as
@@
Note that when implementing the rotation law (55) – that hereafter will refer to as the “CR” law – in the Hydro-RNS code we need to introduce the standard (rescaled) parameters $\hat{j}_{\rm max},\hat{j}_{e},\lambda_{1},\lambda_{2}$, and $A$, but also the new parameters $\beta$ and $\gamma$.
Following the structure of the code with auxiliary $\lambda_{2}$ we are left with five free parameters $\beta$, $\gamma,A,\lambda_{1}$, and $j_{\rm max}/j_{e}$ which, together, control the shape of the profile. In particular, the parameter $\lambda_{1}$ regulates the maximum angular velocity and can be read of from the ratios of the angular velocities from the merger-profiles. The parameter $\beta$ controls instead the slope of the angular velocity for angular momenta smaller than $j_{\rm max}$, which tends to become constant (linear behaviour) for smaller values of $\beta$. The parameters $A$ and $\gamma$ are responsible for the part with angular momenta larger than $j_{\rm max}$, with $\gamma$ controlling how far from the maximum the decreasing part is extended, and $A$ regulating the slope of the angular velocity in the Keplerian region. Note that larger values of the ratio $\gamma/A$ yield profiles that extend to larger angular momenta.
While the CR law provides a more accurate representation of the remnant rotational properties, it also offers additional challenges during the numerical calculation and can easily lead to a loss of convergence in the iterative procedure employed by Hydro-RNS in computing a solution. In particular, for large values of $\lambda_{1}=\Omega_{\rm max}/\Omega_{c}$, the range of possible guesses for $r_{e}$ and of the corresponding solutions of $\Omega_{e}$ grows substantially. In addition, if the guess for $r_{e}$ is rather different from the one for which a solution exists, the code fails to converge because no solution with the guessed $\Omega_{e}$ can be found with the Euler equation. It is in principle possible to manually control the iterations of $r_{e}$ and to extend the range of search of $\Omega_{e}$ so that a solution is found, but this is tedious and not robust. However, an approach that tackles these problems consists “blending” the solution with the new CR rotation law with the solution obtained with the U9 law, which is less accurate but less challenging in the regions of large specific angular momentum. Hence, we introduce a smooth transition from the converging U9 rotation law to the new CR law via a simple linear combination
$\Omega(j)=a\ \Omega_{\rm U9+}(j)+b\ \Omega_{\rm CR}(j)\,,$ (63)
where $a,b\in[0,1]$ and $a+b=1$. Hence, starting with $a=1,\,b=0$ and progressively marching to a situation in which $a=0,\,b=1$, the Hydro-RNS code can find equilibrium solutions with axis ratios that are larger than those allowed by the U9+ rotation law. Clearly, the strategy adopted in the blending (63) with the U9+ law can be employed also with other differential-rotation laws. Figure 11 illustrates the result of this approach by showing how an equilibrium solution with the CR law with $j_{\rm max}/j_{e}=0.82,\beta=0.95,\gamma=1.8,A=20,\lambda_{1}=3$ (red thick line) can be obtained from an initial U9+ law with $p=1,\xi=1,\lambda_{1}=2,\lambda_{2}=1.6$ (blue thick line) via a sequence of intermediate solutions which represent the various flavours of the blending  (63) (blue and red thin lines). The final solution is compared with the angular velocity profile of the DD2 BNS simulation (black thick line) at $t-t_{\rm mer}=20\,{\rm ms}$, showing the superior accuracy of the new CR law. We postpone to Sec. 5.3 the discussion of how the CR-law performs when modelling the remnants of BNS mergers, we here briefly provide evidence that this law allows one to construct numerical models of differentially rotating stars by presenting in Fig. 12 a representative space of solutions for stars modelled with the DD2 EOS and with the following set of fixed parameters: $\hat{j}_{\rm max}/\hat{j}_{e}=0.82$, $\beta=0.95$, $\gamma=1.8$, $A=20$, $\lambda_{1}=3$ and that was obtained after performing a progressive linear blending with the U9+ law. Note that also in this case, the choice of parameters leads to a turning-point stability line such that $\partial M/\partial\rho_{c}|_{\rm tp}>0$. As we comment next, this is not a coincidence and reflects a general behaviour of differential-rotation laws that are inspired by BNS mergers.